1. A bus company increased the bus fare from RM1.50 to RM2.50. After the fare increased, it was reported that the number of passengers taking their buses declined from 100 million to 96

           million.

 

           (i)        Calculate the bus fare price elasticity of demand.                                    (4 marks)

           

           (ii)       Interpret the answer in part (i).                                                                        (2 marks)

           

(iii)      Examine the decision to increase the price of bus fare in terms of its effect on total revenue and profits, assuming costs are unchanged.                                        (3 marks)


1
Expert's answer
2021-02-15T17:49:56-0500

(i) The price elasticity of demand can be calculated as follows:


"E_d=\dfrac{\%\Delta Q}{\%\Delta P},""E_d=\dfrac{\dfrac{Q_2-Q_1}{Q_1+Q_2}}{\dfrac{P_2-P_1}{P_1+P_2}}=\dfrac{\dfrac{96-100}{100+96}}{\dfrac{2.50-1.50}{1.50+2.50}}=-0.081."

(ii) The sign minus means that the demand is inelastic.

(iii) Let's find the TR in case the price doesn't change:


"TR_1=P_1Q_1=100000000\cdot RM1.50=RM150000000."

Now, let's find the TR in case the price increases from RM1.50 to RM2.50:


"TR_2=P_2Q_2=96000000\cdot RM2.50=RM240000000."

Let's find "\Delta TR":


"\Delta TR=TR_2-TR_1=RM240000000-RM150000000=RM90000000."


So, as we can see from the calculations, in case of increasing the bus fare, the bus company earned an additional RM90000000 in revenue.